OpenStudy (anonymous):

15 over 19 t = –9 solve the equation

OpenStudy (anonymous):

is it \[15/(19t)=-9\]or\[(15/19)t=-9\]

OpenStudy (anonymous):

i need it to be explained more of how you came up with it

OpenStudy (anonymous):

The top and bottom equations have different answers, we need to know which one is correct.

OpenStudy (anonymous):

im sorry it says 5 over 13 t=-9

OpenStudy (anonymous):

Are 13 and t both under 5, or is 5 over 13 in front of t?

OpenStudy (anonymous):

5 over 13 is in front of t

OpenStudy (anonymous):

\[(5/13)t=-9\]\[(13/5)(5/13)t=-9(13/5)\]\[t=-117/5\] Cannot be simplified.